What is the work done by a system when it absorbs 50J of heat and gains 24 J of energy?
1
Expert's answer
2021-10-07T03:44:17-0400

"q = 50J\\
∆U = 24J\\
w =\ ?"


"∆U = ∆q - w\\
w= ∆q - ∆U\\
w = 50- 24 = 26 J"

"\therefore" The system expands and the work done is 26J"."